OpenStudy (agervick):

will fan and medal Which is an equation of a circle with center (2, -10) and radius 3? A. (x - 2)2 + (y + 10)2 = 9 B. (x + 10)2 + (y - 2)2 = 9 C. (x - 2)2 + (y + 10)2 = 3 D. (x + 2)2 + (y - 10)2 = 3

OpenStudy (prettygirl_shynice):

okay ready

OpenStudy (agervick):

yes

OpenStudy (prettygirl_shynice):

okay any circle with center (a,b) and radius r has the equation (x-a)^2 + (y-b)^2 = r^2 if you think about it, you will see why this is obvious if you thinking about points on the circle and what its distance from the center should be here (x-2)^2 + (y+1)^2 = 9

OpenStudy (agervick):

Thank you<3

OpenStudy (prettygirl_shynice):

so look at number a and this

OpenStudy (prettygirl_shynice):

ur welcome bye the way text me if u need help im here